OpenStudy (anonymous):

Find an equation of the tangent line to the circle: x^2+y^2=25 at point (-3,4). then graph the line together with the circle.

OpenStudy (anonymous):

in order to graph equations, you need to solve them for y. To graph x^2+y^2=25 plot y=sqrt (25-x^2) and y=-(sort)25-x^2

OpenStudy (anonymous):

I REALLY REALLY NEED QUICK FAST HELP GUYS!

hartnn (hartnn):

to get equation of line, we need slope and a point on that tangent point is givn, do u know how you can get slope ??

hartnn (hartnn):

x^2+y^2= 25 differentiate this w.r.t x, what u get ?

OpenStudy (anonymous):

whats w.r.tx?

OpenStudy (anonymous):

i'm solving for y right now and got y=sqrt(25-x^2)

OpenStudy (anonymous):

now i'm finding the first derivative respect to x right?

hartnn (hartnn):

w.r.t x means with respect to x :P ok, yes for slopee, you'll need dy/dx at x=-3

hartnn (hartnn):

x^2+y^2 is a circle with centre as origin and radius = sqrt 25 = 5

hartnn (hartnn):

**x^2+y^2=25

OpenStudy (anonymous):

right~ ok so far so good!

hartnn (hartnn):

x^2+y^2= 25 differentiate this w.r.t x, what u get ?

OpenStudy (anonymous):

-x/sqrt25-x^2 = y' no? yes?

hartnn (hartnn):

there was no need to isolate 'y' for differentiating..... you could've implicit differentiated x^2+y^2 = 25 to get 2x + 2y (dy/dx) = 0 got this ?? now isolate dy/dx from that

hartnn (hartnn):

derivative of y^2 will be 2y (dy/dx) because of chain rule

OpenStudy (anonymous):

so what I did was wrong?

hartnn (hartnn):

not incorrect, but not required...

hartnn (hartnn):

you'll get same slope with what you have done and inplicit differentiation... why not take an easier way when you have it :)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so then its -2x/2y right?

hartnn (hartnn):

yes! dy/dx = -x/y and slope of tangent will be dy/dx at x =-3, y =4 so, slope = m = -(-3)/4 = ... ? getting this ?

OpenStudy (anonymous):

yup so far!

hartnn (hartnn):

now you have slope and a point also on the tangent can you find the equation ??

hartnn (hartnn):

use, y-y1 = m (x-x1)

OpenStudy (anonymous):

right but how do I plot the circle when the plots are like y= sqrt 25-x^2?

hartnn (hartnn):

as i said, \(x^2+y^2 =r^2 \) is a circle with origin as centre (0,0) and radius = r here, r=5

OpenStudy (anonymous):

why are they giving me two y's?

OpenStudy (anonymous):

ok so what if I know the radius? what does that imply?

hartnn (hartnn):

to plot a circle, you only need centre and radius! like circle with radius = r and centre as origin will be plot as